What is the resultant temperature when 100g of steam at 100°C is passed through 500 g of ice at −20°C. The specific heat

of ice is 0.5 cal / g°C.


1
Expert's answer
2021-05-03T10:31:26-0400

Q = mcΔT

Q = mL

Q(absorbed by ice) = Q(released by stem)

"Q_1 + Q_2 = Q_3 + Q_4 \\

m_{ice}L_f + m_{ice}c_i(T_f – T_i) = m_{steam}L_v + m_{steam}c_w(T_f -T_i) \\

L_f = 79.72 \; cal/g \\

L_v = 540 \; cal/g \\

c_w = 1 \; cal/gC"

"500 \times 79.72 + 500 \times 0.5 (T_f - (-20)) = 100 \times 540 + 100 \times 1 (T_f -100) \\

39860 + 250T_f + 5000 = 54000 + 100T_f -10000 \\

150T_f = -860 \\

T_f = -5.73 \; °C"